OpenStudy (anonymous):

Write the equation of the line that passes through (2, 3) and is parallel to the line 12x – 5y = 2.

OpenStudy (a1234):

y - y1 = m(x - x1) y - 3 = -5(x - 2) y - 3 = -5x + 10 y = -5x + 13

OpenStudy (anonymous):

y = 12/5x - 9/5 y = 5/12x - 9/5 y = 5/2 x + 9/5 y = 5/12x - 9/5

OpenStudy (anonymous):

@a1234

OpenStudy (a1234):

I'm so sorry...I typed in a rush. Just a sec...

OpenStudy (a1234):

12x - 5y = 2 -5y = -12x + 2 y = 12/5x - 2/5 y - y1 = m(x - x1) y - 3 = 12/5(x - 2) y - 3 = 12/5x - 24/5 y = 12/5x - 1.8 So it is A.

OpenStudy (anonymous):

Can you help me with one more?